OpenStudy (anonymous):

If y varies directly as x and y is 6 when x is 72, what is the value of y when x is 8?

hero (hero):

The formula for direct variation is y = kx In the first case, when y = 6 and x = 72: k = y/x = 6/72 = 6/(6*12) = 1/12 In the second case, when x = 8: y =(8)(1/12) = 8/12 = 2/3 so y = 2/3 when x = 8

OpenStudy (anonymous):

fast reply hero

hero (hero):

I don't think it was fast enough.